### v5

@@ 5.1 Which Global Variables Correlate with our Principal Components?
(1) (2) (3) (4) (5) (6)
PC1 PC2 PC3
VIX -0.440 -0.489 0.010 0.025 0.145 -0.176
(0.153)*** (0.171)*** (0.135) (0.126) (0.171) (0.184)
Oil Price -0.406 -0.337 -0.145 -0.129 0.217 0.417
(0.222)* (0.227) (0.191) (0.192) (0.224) (0.211)*
Shadow Rate 1.092 1.041 1.167 1.166 1.077 0.853
(0.098)*** (0.101)*** (0.096)*** (0.094)*** (0.108)*** (0.115)***
Non-Bank Share -0.783 -0.608 0.642 0.682 -1.048 -0.550
(0.236)*** (0.246)** (0.224)*** (0.236)*** (0.279)*** (0.261)**
World GDP Growth 0.215 0.201 -0.269 -0.272 0.183 0.144
(0.125)* (0.116)* (0.086)*** (0.081)*** (0.134) (0.112)
World Macropru. -2.951 -3.016 -0.586 -0.587 1.713 1.428
(0.205)*** (0.209)*** (0.190)*** (0.194)*** (0.210)*** (0.186)***
World Inflation 0.063 0.026 0.347 0.339 -0.003 -0.111
(0.126) (0.130) (0.080)*** (0.082)*** (0.108) (0.106)
US REER -0.784 -0.647 0.802 0.824 -0.693 -0.241
(0.257)*** (0.256)** (0.241)*** (0.232)*** (0.288)** (0.284)
GFC 0.563 0.065 2.043
(0.290)* (0.235) (0.456)***
COVID -0.333 -0.230 0.120
(0.625) (0.382) (0.579)
Time Periods 72 72 72 72 72 72
$R^{2}$ 0.971 0.973 0.920 0.921 0.675 0.786
@@
Regressions of the PCs on observed common factors, with PCs obtained from Table 2 and all regressors being standardized. Global Financial Crisis (GFC) and COVID are dummies equal to one for 2007:Q4-2011:Q4, following Laeven and Valencia (2018), and 2019:Q4-2020:Q4, respectively. All variables are defined as in Table 9.
@@
We address this concern here by relating the PCs to observable common factors that have been featured in previous studies and also in the earlier specifications in Tables 2. This is done in Table 6, which reports regressions of each PC on that set of observed global factors, with the latter being standardized by subtracting the mean and dividing by the standard deviation of each variable.
@@
As becomes apparent, the first PC loads by far most significantly on the macropudential stringency variable (column 1). In particular, tighter macroprudential regulations reduce banks’ non-core ratios. The same is true for a higher VIX, higher oil prices, a higher world non-bank share, a lower US shadow rate and a stronger USD, but their individual effects on the PC and hence banks’ non-core ratios are smaller. Column (3) of Table 6 shows that PC2 is mostly driven by the US shadow rate, with a correlation between both variables being positive. At first sight, this positive relation between US rates and banks’ foreign borrowing seems inconsistent with the previous global financial cycle literature (e.g., Miranda-Agrippino and Rey, 2020), which shows that lower US rates increase bank leverage and cross-border capital flows. Note, however, that changes in the US shadow rate during our sample period are predominantly driven by the Fed’s quantitative easing and tightening cycles. Larger central bank purchases of long-term bonds (lower shadow rates) reduce long-term yields and flatten the yield curve (D’Amico and King, 2013; Sudo and Tanaka, 2021). This, in turn, should then incentivize banks to fund any lending expansions with long-term bond issuances and not with short-term borrowing via the interbank market. As we show in Section 5.2, the changes in non-core ratios we observe are mostly driven by changes in interbank borrowing, so it is not a surprise to see that lower US shadow rates tend to reduce banks’ non-core ratios. In fact, when we use the loan-to-deposit ratio as outcome variable and correlate the PCs extracted from this specification with the US shadow rate, we obtain the same negative coefficient as the extant literature (see Appendix Table 11). The same is true when we correlate the US shadow rate and PC2 extracted from our non-core ratio for a sample period that excludes any quantitative easing or tightening policies. For instance, before 2007, the correlation between both variables is -16.3%.
@@
Finally, column (5) of Table 6 shows that PC3 also correlates most significantly with macroprudential policies, but the coefficient estimate is smaller than that in column (1) and the $R^{2}$ of this regression is quite small (0.68). A graphical inspection of this PC in Figure 3 shows that PC3 tends to increase during crises. We hence add to the previous specification two crisis dummies—one US banking crisis indicator that is one between 2007:Q4 and 2011:Q4222222Laeven and Valencia (2018) identify a US banking crises for this period and one Covid dummy that takes a value of one for the initial phase of the Covid pandemic, i.e., between 2019:Q4 and 2020:Q4. For the sake of completeness, we also add the same crisis dummies to the regressions of PC1 and PC2. As columns (2), (4) and (6) show, adding the crisis dummies significantly increases the $R^{2}$ for the PC3 regression (column 6). The US banking crisis dummy further turns out being economically most significant and clearly dominates the macroprudential stance variable. In contrast, the crisis dummies do not affect PC1 and PC2 to the same extent as PC3.
@@
We have shown before that the loadings to PC1 and PC2 are stronger for countries with a fixed exchange rate, whereas the effect of PC3 is quite similar in economic terms for fixers and floaters. Therefore, having a floating exchange rate helps insulating against influences stemming from regulatory, financial or real variables, all of which are important drivers of PC1 and PC2, whereas it does not help insulating against the effects of financial crises, which seem to drive PC3. This novel result is consistent with column (4) of Table 4.
@@ 5.2 A Breakdown of Banks’ Foreign Liabilities
(1) (2) (3) (4) (5) (6)
Total Foreign Domestic Banks Non-Banks Other
Lag Dep. Var. 0.958 0.957 0.950 0.954 0.961 0.951
(0.015)*** (0.015)*** (0.018)*** (0.019)*** (0.013)*** (0.015)***
PC1 0.196 0.100 0.109 0.198 0.017 -0.050
(0.067)*** (0.032)*** (0.043)** (0.060)*** (0.017) (0.023)**
PC2 0.250 0.106 0.140 0.218 0.049 -0.106
(0.127)** (0.058)* (0.089) (0.111)** (0.039) (0.041)***
PC3 0.400 0.130 0.282 0.392 0.015 0.042
(0.221)* (0.096) (0.147)* (0.175)** (0.058) (0.061)
Obs 1633 1627 1627 1554 1633 1554
$R^{2}$ 0.981 0.981 0.977 0.976 0.981 0.963
@@
Table 7: BIS Variables Regressed on Principal Components
Table 7 contains the attendant results which show that all three principal components are positively correlated with banks’ total cross-border BIS liabilities, consistent with our benchmark results (column 1). The breakdown into the different components, shown in columns (2)-(6), indicate that this increase materializes via both increases in domestic- and foreign-currency liabilities, as well as an increase in banks’ liabilities vis-a-vis banks. In contrast, liabilities vis-a-vis other financials or all other sectors are hardly (in some cases even negatively) correlated with our principal components. As Table 12 of the Appendix shows, most of these effects are again stronger for countries on a fixed exchange rate. Overall, this is evidence that the relation between the global financial cycle and non-core ratios documented in this paper works via an increase in banks’ foreign funding provided via the interbank market—a funding source that is well-known to be quite flighty. The currency denomination does not seem to matter a lot.
@@ 6 Conclusion
A key finding is that flexible exchange rates play a significant insulation role in the way global factors affect the ratio of non-core to core funding. In particular, global shocks affect non-core ratios less for countries on a flexible exchange rate regime and even more strongly outside of global crises episodes, like the GFC or the Covid pandemic. We also document that tighter macroprudential regulations help insulate from the global financial cycle.
@@ Appendix A Data Appendix
In this appendix, we describe in detail how we constructed our country-level panel data set. Table 8 contains the countries included in the analysis, while Table 9 shows an overview of the variables employed in the analysis.
@@
Fixed: =1 for countries with strictly fixed peg, i.e., countries where the coarse Ilzetzki et al. (2019) exchange rate classification is equal to 1. The data come at monthly frequency and we use end-of-quarter values for our quarterly data. The variable coverage ends in 2019, so we assume that countries do not switch their classification after that. [^23] For our sample, the Czech Republic is the only country that switches its exchange rate regime (from floating to fixed in 2014). We define the country as a floater throughout, but results are similar if we don’t do so.
@@
VIX: The CBOE Volatility index. Source: Fred.
@@
Country Fixed Exchange Rate Macro Prudential Tightness
Australia 0 0
Austria 1 1
Belgium 1 0
Canada 0 1
Chile 0 0
Cyprus 1 0
Czech Republic 0 0
Denmark 1 1
Estonia 1 0
Finland 1 1
France 1 0
Germany 1 0
Greece 1 0
Hungary 0 1
Iceland 0 1
Ireland 1 0
Israel 0 1
Italy 1 0
Japan 0 0
Korea, Republic of 0 1
Luxembourg 1 1
Netherlands 1 1
Norway 0 1
Poland 0 1
Portugal 1 1
Romania 0 1
Slovenia 1 0
Spain 1 0
Sweden 0 1
United Kingdom 0 0
United States 0 0
@@ Appendix B Additional Tables and Figures
non-core 3 3
loan to deposit 5 4
Obs. CF 7 7
@@
Number of factors estimations based on Growth Rate (GR) estimator from Ahn and Horenstein (2013). Variables in Observed common factors (Obs. CF) are: VIX, Oil Price, Shadow Rate, Non-Bank Share, World GDP Growth, World Inflation, World Macroprudential and the US real exchange rate. Estimations are done in Stata using xtnumfac (Ditzen and Reese, 2023).
@@
(1) (2) (3) (4) (5) (6)
PC1 PC2 PC3
VIX -0.601 -0.663 0.220 0.015 1.037 0.883
(0.201)*** (0.259)** (0.228) (0.238) (0.196)*** (0.229)***
Oil Price -0.571 -0.489 -0.418 -0.268 0.340 0.407
(0.254)** (0.275)* (0.317) (0.333) (0.294) (0.296)
Shadow Rate 1.105 1.043 -0.541 -0.693 0.235 0.140
(0.148)*** (0.164)*** (0.158)*** (0.195)*** (0.168) (0.181)
Non-Bank Share -0.844 -0.636 -1.025 -0.650 -0.142 0.024
(0.258)*** (0.268)** (0.329)*** (0.333)* (0.283) (0.310)
World GDP Growth 0.183 0.166 0.299 0.269 -0.081 -0.094
(0.163) (0.155) (0.164)* (0.151)* (0.140) (0.134)
World Macropru. -2.547 -2.627 1.612 1.418 -0.505 -0.626
(0.261)*** (0.279)*** (0.273)*** (0.288)*** (0.273)* (0.287)**
World Inflation -0.129 -0.174 -0.167 -0.248 0.585 0.549
(0.153) (0.161) (0.149) (0.168) (0.153)*** (0.156)***
US REER -1.169 -1.005 -1.730 -1.402 1.014 1.182
(0.320)*** (0.358)*** (0.427)*** (0.472)*** (0.399)** (0.409)***
GFC 0.678 1.447 0.796
(0.421) (0.538)*** (0.382)**
COVID -0.373 -0.128 0.313
(0.824) (0.609) (0.638)
Time Periods 72 72 72 72 72 72
$R^{2}$ 0.946 0.948 0.756 0.782 0.634 0.651
@@
Regressions of the PCs extracted from the loan-to-deposit regression in Table 4 on the observed common factors. Variables are standardized. Global Financial Crisis (GFC) and COVID are dummies equal to one for 2007:Q4-2011:Q4, following Laeven and Valencia (2018), and 2019:Q4-2020:Q4, respectively. All variables are defined as in Table 9.
(1) (2) (3) (4) (5) (6)
Total Foreign Domestic Banks Non-Banks Other
Lag Dep. Var. 0.954 0.952 0.948 0.949 0.960 0.950
(0.016)*** (0.015)*** (0.018)*** (0.021)*** (0.013)*** (0.016)***
Principal Components (fixed = 0)
PC1 0.052 0.033 0.020 0.110 -0.007 -0.049
(0.055) (0.046) (0.017) (0.038)*** (0.014) (0.031)
PC2 0.003 -0.016 0.024 0.034 0.026 -0.058
(0.084) (0.067) (0.033) (0.073) (0.025) (0.031)*
PC3 0.259 0.239 0.036 0.309 0.031 -0.075
(0.147)* (0.118)** (0.049) (0.114)*** (0.041) (0.058)
Principal Components (fixed = 1)
PC1 0.257 0.121 0.155 0.195 0.041 -0.010
(0.124)** (0.065)* (0.072)** (0.100)* (0.031) (0.042)
PC2 0.410 0.207 0.191 0.360 0.040 -0.096
(0.218)* (0.109)* (0.150) (0.221) (0.068) (0.077)
PC3 0.243 -0.174 0.410 0.118 -0.026 0.209
(0.372) (0.180) (0.243)* (0.313) (0.102) (0.118)*
Obs 1633 1627 1627 1554 1633 1554
$R^{2}$ 0.981 0.981 0.977 0.976 0.981 0.963
@@
Table 12: BIS Variables Regressed on Principal Components:
The Role of a Fixed Exchange Rate

### v6

@@ 5.1 Which Global Variables Correlate with our Principal Components?
Regressions of the PCs on observed common factors, with PCs obtained from Table 4.1 and all regressors being standardized. Global Financial Crisis (GFC) and COVID are dummies equal to one for 2007:Q4-2011:Q4, following Laeven and Valencia (2018), and 2019:Q4-2020:Q4, respectively. All variables are defined as in Table 9.
@@
We address this concern here by relating the PCs to observable common factors that have been featured in previous studies and also in the earlier specifications in Tables 4.1. This is done in Table 5.1, which reports regressions of each PC on that set of observed global factors, with the latter being standardized by subtracting the mean and dividing by the standard deviation of each variable.
@@
As becomes apparent, the first PC loads by far most significantly on the macropudential stringency variable (column 1). In particular, tighter macroprudential regulations reduce banks’ non-core ratios. The same is true for a higher VIX, higher oil prices, a higher world non-bank share, a lower US shadow rate and a stronger USD, but their individual effects on the PC and hence banks’ non-core ratios are smaller. Column (3) of Table 5.1 shows that PC2 is mostly driven by the US shadow rate, with a correlation between both variables being positive. At first sight, this positive relation between US rates and banks’ foreign borrowing seems inconsistent with the previous global financial cycle literature (e.g., Miranda-Agrippino and Rey, 2020), which shows that lower US rates increase bank leverage and cross-border capital flows. Note, however, that changes in the US shadow rate during our sample period are predominantly driven by the Fed’s quantitative easing and tightening cycles. Larger central bank purchases of long-term bonds (lower shadow rates) reduce long-term yields and flatten the yield curve (D’Amico and King, 2013; Sudo and Tanaka, 2021). This, in turn, should then incentivize banks to fund any lending expansions with long-term bond issuances and not with short-term borrowing via the interbank market. As we show in Section 5.2, the changes in non-core ratios we observe are mostly driven by changes in interbank borrowing, so it is not a surprise to see that lower US shadow rates tend to reduce banks’ non-core ratios. In fact, when we use the loan-to-deposit ratio as outcome variable and correlate the PCs extracted from this specification with the US shadow rate, we obtain the same negative coefficient as the extant literature (see Appendix Table B). The same is true when we correlate the US shadow rate and PC2 extracted from our non-core ratio for a sample period that excludes any quantitative easing or tightening policies. For instance, before 2007, the correlation between both variables is -15.7%.
@@
Finally, column (5) of Table 5.1 shows that PC3 also correlates most significantly with macroprudential policies, but the coefficient estimate is smaller than that in column (1) and the $R^{2}$ of this regression is quite small (0.68). A graphical inspection of this PC in Figure 3 shows that PC3 tends to increase during crises. We hence add to the previous specification two crisis dummies—one US banking crisis indicator that is one between 2007:Q4 and 2011:Q4232323Laeven and Valencia (2018) identify a US banking crisis for this period. and one Covid dummy that takes a value of one for the initial phase of the Covid pandemic, i.e., between 2019:Q4 and 2020:Q4. For the sake of completeness, we also add the same crisis dummies to the regressions of PC1 and PC2. Adding the crisis dummies significantly increases the $R^{2}$ for the PC3 regression (column 6). The US banking crisis dummy further turns out being economically most significant and clearly dominates the macroprudential stance variable. In contrast, the crisis dummies do not affect PC1 and PC2 to the same extent as PC3 (columns 2 and 4).
@@
We have shown before that the loadings to PC1 and PC2 are stronger for countries with a fixed exchange rate, whereas the effect of PC3 is quite similar in economic terms for fixers and floaters. Therefore, having a flexible exchange rate helps insulating against influences stemming from regulatory, financial or real variables, all of which are important drivers of PC1 and PC2, whereas it does not help insulating against the effects of financial crises, which seem to drive PC3. This novel result is consistent with column (4) of Table 4.2.
@@ 5.2 A Breakdown of Banks’ Foreign Liabilities
Table 7: BIS Variables Regressed on Principal Components
@@
Table 5.2 contains the attendant results which show that all three principal components are positively correlated with banks’ total cross-border BIS liabilities, consistent with our benchmark results (column 1). The breakdown into the different components, shown in columns (2)-(6), indicate that this increase materializes via both increases in domestic- and foreign-currency liabilities, as well as an increase in banks’ liabilities vis-a-vis banks. In contrast, liabilities vis-a-vis other financials or all other sectors are hardly (in some cases even negatively) correlated with our principal components. As Table B of the Appendix shows, at least some of these effects are again stronger for countries on a fixed exchange rate. Overall, this is evidence that the relation between the global financial cycle and non-core ratios documented in this paper works via an increase in banks’ foreign funding provided via the interbank market—a funding source that is well-known to be quite flighty. The currency denomination does not seem to matter a lot in our sample. [^24] This may or may not hold for a broader sample that includes emerging markets where currency mismatches between banks’ assets and liabilities are much larger and exchange rates are far more volatile than in our sample of advanced countries. So, we acknowledge that this result may be somewhat sample specific.
@@ 6 Conclusion
A key finding is that flexible exchange rates play a significant insulation role in the way global factors affect the ratio of non-core to core funding. In particular, global shocks affect non-core ratios less for countries on a flexible exchange rate regime and even more strongly outside of global crises episodes, like the GFC or the Covid pandemic. We also document that tighter macroprudential regulations help insulate bank funding from the global financial cycle.
@@ Appendix A Data Appendix
In this appendix, we describe in detail how we constructed our country-level panel data set. Table A contains the countries included in the analysis, while Table 9 shows an overview of the variables employed in the analysis.
@@
Fixed: =1 for countries with strictly fixed peg, i.e., countries where the coarse Ilzetzki et al. (2019) exchange rate classification is equal to 1. The data come at monthly frequency and we use end-of-quarter values for our quarterly data. The variable coverage ends in 2019, so we assume that countries do not switch their classification after that. [^25] For our sample, the Czech Republic is the only country that switches its exchange rate regime (from flexible to fixed in 2014). We define the country as a floater throughout, but results are similar if we don’t do so.
@@
VIX: The log of the CBOE Volatility index. Source: Fred.
@@ Appendix B Additional Tables and Figures
Non-Core Ratio 3 3
Loan-to-Deposit Ratio 5 4
Observed Common Factor 7 7
@@
Number of factors estimations based on Growth Rate (GR) estimator from Ahn and Horenstein (2013). Variables in Observed common factors are: VIX, Oil Price, Shadow Rate, Non-Bank Share, World GDP Growth, World Inflation, World Macroprudential and the US real exchange rate. Estimations are done in Stata using xtnumfac (Ditzen and Reese, 2023).
@@
Regressions of the PCs extracted from the loan-to-deposit regression in Table 4.2 on the observed common factors. Variables are standardized. Global Financial Crisis (GFC) and COVID are dummies equal to one for 2007:Q4-2011:Q4, following Laeven and Valencia (2018), and 2019:Q4-2020:Q4, respectively. All variables are defined as in Table 9.
Table 12: BIS Variables Regressed on Principal Components: The Role of a Fixed Exchange Rate